OpenStudy (anonymous):

pairs of functions, four problems in one 46 tiawanad f(x)=3x g(x)=7x^2-4x+6 a) 3(7x^2-4x+6) 21x^2-12x+18

OpenStudy (anonymous):

then you have to do (fog)(-3) 21(-3)^2-12(-3)+18 63^2-36+18

OpenStudy (anonymous):

or would you do 3*3=9*21=189